OpenStudy (anonymous):

Could someone explain how to set this problem up as an algebraic equation: The sides of a triangle are in the ratio 5:12:13. What is the length of each side of the triangle if the perimeter of the triangle is 15 inches?

OpenStudy (anonymous):

5x ,12x and 13x are the sides of the triangle Perimeter of triangle is sum of sides of triangle 5x +12x +13x = 15 inches 30x = 15 inches x = 15/30 = 1/2 = 0.5 inch 5x = 5*0.5 = 2.5 inches 13x = 13*0.5 = 6.5 inchers 12x = 12*0.5 = 6 inches

OpenStudy (amistre64):

5:12:13, P=30 : /2 5/2 : 6 :13/2, P=15

OpenStudy (amistre64):

since the ration never changes; all the measures have to be scaled together to keep it the same; like fractions

OpenStudy (anonymous):

I did the check of your answer Rita: (Because I'm absolutely clueless with word problems and have to learn them step by step.) 5x = 5*0.5 = 2.5 13x = 13*0.5 = 6.5 12x = 12*0.5 = 6 2.5+6.5+6=15 It adds up. Thanks. =) Someone tried to get me to do it the way you did Amistre, but I think it's easier for me to do it the way Rita suggested. Thanks a ton though for the reply. =)

OpenStudy (amistre64):

its all good ;)